# Sampling Theory for Function Approximation with Numerical Redundancy

### v1

@@ 3 Discrete regularized least squares approximations
The close relationship between numerically redundant spanning sets and $\ell^{2}$-regularization motivates the study of discrete regularized least squares approximations, which is detailed in this section. We begin by analysing regularized least squares approximations with a more general penalty term, as the methodology we use naturally extends to such a setting. In section 3.2, we discuss $\ell^{2}$-regularization, particularly for function approximation with a numerically redundant spanning set.
3.1 Regularized approximation with a general penalty term
@@
with associated operator $\mathcal{F}_{n,m}^{\text{RLS}}$ defined by (7). We assume that $\mathcal{L}_{n,m}^{\text{RLS}}$ is well defined, i.e., that the minimizer is unique. The penalty function $p_{\epsilon}:\mathbb{C}^{n^{\prime}}\to\mathbb{R}^{+}$ can depend on a regularization parameter $\epsilon$. Furthermore, we assume that $p_{\epsilon}$ is even and that $\sqrt{p_{\epsilon}}$ is subadditive, meaning that for all $\mathbf{x},\mathbf{y}\in\mathbb{C}^{n^{\prime}}$ it holds that
@@
$\leq$ $\|\mathcal{T}_{n}\mathbf{x}-f\|_{m}+\|\mathbf{f}-\mathcal{M}_{m}\mathcal{F}_{n,m}^{\text{RLS}}\mathbf{f}\|_{2}+\|\mathbf{n}\|_{2}+\sqrt{p_{\epsilon}(\mathbf{x})}+\sqrt{p_{\epsilon}(\mathcal{L}_{n,m}^{\text{RLS}}\mathbf{f})},$
$\leq$ $(1+\sqrt{2})(\|\mathcal{T}_{n}\mathbf{x}-f\|_{m}+\sqrt{p_{\epsilon}(\mathbf{x})}+\|\mathbf{n}\|_{2}),$
@@
3.2 Discretizing numerically redundant spanning sets
We are specifically interested in $\ell^{2}$-regularization in order to compute accurate approximations in a numerically redundant spanning set. In this case, one has $p_{\epsilon}(\mathbf{x})=\epsilon^{2}\|\mathbf{x}\|_{2}^{2}$, such that the discretization condition becomes
@@ 3.2.2 Interpretation of the discretization condition
and, hence, for these sets of expansion coefficients (31) is automatically satisfied for $A^{\epsilon}_{n,m}=1$. More generally, (31) is automatically satisfied for any $\mathbf{x}\in\mathbb{C}^{n^{\prime}}$ satisfying $\|\mathcal{T}_{n}\mathbf{x}\|_{H}\leq\epsilon\|\mathbf{x}\|_{2}$. This reflects the idea that regularization makes the effective approximation space smaller. What happens in the case of an orthonormal spanning set? Such a set satisfies Parseval’s identity: $\|\mathcal{T}_{n}\mathbf{x}\|_{H}=\|\mathbf{x}\|_{2}$. Therefore, (31) is equivalent to
@@
In order to further illustrate the effect of the penalty term on the discretization condition, we present a simple example. Suppose one has a basis $\Phi_{n}$ and an associated sampling operator $\mathcal{M}_{m}$, and the basis is augmented with one extra function $\psi$. Typically, to ensure a good discretization, you would also need to enhance $\mathcal{M}_{m}$. However, if $\psi$ can be approximated to machine precision within the span of $\Phi_{n}$, it is intuitively clear that adding this extra function does not increase the approximation space “numerically”. In other words, the increase in approximation space obtained by adding an extra function to the spanning set is in such cases invisible on a computer with finite precision. This phenomenon is also discussed by Boyd in [15, section 16.6]. Consequently, it is also reasonable to expect that no additional samples are required. This intuition can be substantiated using (31), which takes into account the effects of finite precision. In general, (9) is not satisfied for the augmented basis, i.e., $A_{n,m}=0$.
@@ 3.2.3 Common methods for $\boldsymbol{\ell^{2}}$-regularization
For completeness, we show that TSVD regularization indeed satisfies (36). The same result follows from combining [20, Lemma 3.3] and [6, Theorem 3.8].
@@ 4 Random sampling for $\boldsymbol{L^{2}}$-approximations
The question naturally arises: how can one construct a sampling operator that satisfies the regularized norm inequality (31)? To address this question, we draw inspiration from the unregularized case. Recently, Cohen and Migliorati [19] introduced an elegant strategy to generate random pointwise samples that satisfy (9) for least squares fitting. This approach, reviewed in section 4.1, is closely related to the concept of leverage scores, which are used to subsample tall-and-skinny matrices in a fully discrete setting. Leverage score sampling is a well-established technique in numerical linear algebra and machine learning. Moreover, an extension of leverage scores, known as ridge leverage scores, has been developed specifically for regularized settings. Using a continuous analogue of these scores, we present a sampling strategy for $\ell^{2}$-regularized function approximation in section 4.2. In section 4.3, a general framework is given for the analysis of the effects of regularization given a numerically redundant spanning set. Throughout the whole section, we work under the mild assumption that for any x in the domain of $V_{n}$, there exists a $v\in V_{n}$ such that $v(\mathbf{x})\neq 0$.
@@ 4.1 Results without regularization
where $\preceq$ denotes the Loewner order. Here, $G_{n,m}$ and $G_{n}$ are the discrete and continuous Gram matrices representing the operators $(\mathcal{M}_{m}\mathcal{U}_{n})^{*}(\mathcal{M}_{m}\mathcal{U}_{n})$ and $\mathcal{U}_{n}^{*}\mathcal{U}_{n}$, respectively, where $\mathcal{U}_{n}$ is the synthesis operator associated with any set of functions $\Psi_{n}$ satisfying $\operatorname{span}(\Psi_{n})=V_{n}$. We stress that this equivalence holds for any such set $\Psi_{n}$, not only for the chosen spanning set $\Phi_{n}$ (1). However, if one considers an orthonormal set $\Psi_{n}=\{\psi_{i}\}_{i=1}^{n}$, the analysis simplifies significantly. In this case, one has $G_{n}=I$ and $A_{n,m}$ is simply related to the smallest eigenvalue of $G_{n,m}$.
As explained in [19, Proof of Theorem 2.1], suppose one uses a sampling operator $\mathcal{M}_{m}$ based on random pointwise samples
@@
This definition assumes that $\Psi_{n}$ is orthonormal. However, it is independent of the specific choice of the orthonormal set $\Psi_{n}$; the function $k_{n}$ only depends on $V_{n}=\operatorname{span}(\Psi_{n})$ and the probability density $\rho$. The function $k_{n}$ has been analysed in the field of classical analysis long before its introduction in the analysis of discrete least squares problems. In classical analysis, it is known as the inverse of the Christoffel function [42]. Furthermore, its discrete analogue is known as leverage scores, which are widely used in randomized numerical linear algebra and machine learning, see e.g. [39, §2].
This reasoning led to the following sampling result, which is a restatement of [23, Lemma 2.1].
@@
[23, Lemma 2.1] Consider the sampling operator $\mathcal{M}_{m}$ defined by the sampling functionals (39). Let $\gamma>0$ and
@@
It follows that the sampling distribution (42) requires only $\mathcal{O}(n\log(n))$ samples for accurate least squares fitting with high probability. As a result, the amount of required data is nearly optimal, where “optimal” corresponds to the case of interpolation with $m=n$. For a summary of optimality benchmarks and probabilistic error analyses, we refer to [23, section 1 and 2]. Also, we note that considerable efforts have been made to obtain $\mathcal{O}(n)$ sampling strategies, see e.g. [1, section 8] and references therein.
4.2 Introducing $\boldsymbol{\ell^{2}}$-regularization
@@
An important difference compared to section 4.1 is the dependence on the chosen spanning set $\Phi_{n}$. In this case, we are bound to the choice $\Psi_{n}=\Phi_{n}$, i.e., $G_{n}$ and $G_{n,m}$ represent the operators $(\mathcal{M}_{m}\mathcal{T}_{n})^{*}(\mathcal{M}_{m}\mathcal{T}_{n})$ and $\mathcal{T}_{n}^{*}\mathcal{T}_{n}$, respectively, where $\mathcal{T}_{n}$ is the synthesis operator associated with the spanning set $\Phi_{n}$.
Similar spectral bounds have been analysed before in the setting of randomized algorithms [9, 10, 11, 44] using analogous techniques as the ones described in section 4.1. In this case, a continuous analogue of the $\epsilon$-ridge leverage scores [9, Definition 1] is of interest
@@
In the discrete context, the quantity $n^{\epsilon}$ is referred to as the effective dimension [9]. It satisfies $n^{\epsilon}\leq n$ and it is a decreasing function of the regularization parameter $\epsilon$.
Using very similar techniques as [10, proof of Lemma 6], we can prove the following sampling result. The reworked proof can be found in Appendix B.
@@
$m\geq\frac{32}{3}\|wk_{n}^{\epsilon}\|_{L^{\infty}(X)}\log(16n^{\epsilon}/\gamma),$
@@
It is worthwhile to interpret the differences between the results laid out in section 4.1 and those in this section. Using Theorem 1, it follows that the sampling strategy proposed in Corollary 1 results in a least squares approximation with
@@
samples, again with high probability and assuming the data $\mathbf{f}$ is noiseless. In these statements the constant $C>0$ is generic and may change with each occurrence. These results demonstrate that regularization generally decreases the achievable accuracy, yet also decreases the amount of data required to obtain a near-best approximation. Furthermore, the effective dimension $n^{\epsilon}$ in regularized approximations serves a similar role as the dimension $n$ in unregularized approximations.
Two important remarks are to be made. First, as discussed in section 2, optimality in the sense of (47) is unattainable when computing with a numerically redundant spanning set on a machine with finite precision. Additionally, section 6 demonstrates that straightforward numerical orthogonalization of the spanning set does not mitigate the effects of finite precision. As a result, optimality in the sense of (49) is the best one can aim for in this case. Second, in practice one typically has $n^{\epsilon}=\mathcal{O}(n)$ such that (48) and (50) are asymptotically equivalent as $n\to\infty$. This suggests that the gain in data efficiency achieved through regularization is generally limited to a constant factor when samples are drawn from the optimal distributions proposed in Corollary 1 and 2. However, one cannot always choose which distribution to sample from. Since $k_{n}$ can differ significantly from $k_{n}^{\epsilon}$, the required number of samples drawn from a specific distribution $\mu$ may vary considerably depending on whether or not the effects of regularization are accounted for, as described in Theorems 9 and 10. This phenomenon is illustrated further on in section 5.2.2 for a Fourier extension frame.
@@ 4.3 Behaviour of $\boldsymbol{n^{\epsilon}}$ and $\boldsymbol{k_{n}^{\epsilon}}$
For approximation with numerically redundant spanning sets, we are specifically interested in the setting where $\epsilon$ is of the order of a finite working precision $\operatorname{\epsilon_{\text{mach}}}$ associated with the numerical computations. It is natural to wonder whether this has a big influence, i.e., whether the effective degrees of freedom $n^{\epsilon}$ differ significantly from $n$ and the behaviour of $k^{\epsilon}_{n}$ differs significantly from the behaviour of $k_{n}$, for small regularization parameters $\epsilon$.
@@
The region in between $i_{1}$ and $i_{2}$ is often referred to as the plunge region [21]. These regions have a different influence on the effective degrees of freedom $n^{\epsilon}$ and the maximum of the inverse effective Christoffel function $k_{n}^{\epsilon}$, as quantified in the following theorem.
@@
where $\operatorname{S_{\text{tail}}}=\sum_{i=i_{2}+1}^{n}\left(\sigma_{i}/\epsilon\right)^{2}$ and $k^{\text{trunc}}_{n}(\mathbf{x})=\sum_{i=1}^{i_{2}}\lvert u_{i}(\mathbf{x})\rvert^{2}$. Furthermore, if $\Phi_{n}$ is linearly independent and uniformly bounded, $\|\phi_{i,n}\|_{L^{\infty}(X)}\leq 1\;(1\leq i\leq n)$, then
$\|k_{n}^{\epsilon}\|_{L^{\infty}(X)}\leq\frac{n}{1-\epsilon^{2}}+((i_{2}-i_{1})+\operatorname{S_{\text{tail}}})\max_{i_{1}<i\leq n}\|u_{i}\|_{L^{\infty}(X)}^{2}.$ (55)
Proof.
For (54), consider
and
In order to obtain (55), we split the above as
$\|k_{n}^{\epsilon}\|_{L^{\infty}(X)}$ $\leq\sup_{\mathbf{x}\in X}\left(\sum_{i=1}^{i_{1}}\lvert u_{i}(\mathbf{x})\rvert^{2}+\sum_{i=i_{1}+1}^{i_{2}}\lvert u_{i}(\mathbf{x})\rvert^{2}+\operatorname{S_{\text{tail}}}\max_{i_{2}<i\leq n}\lvert u_{i}(\mathbf{x})\rvert^{2}\right)$
$\leq\sup_{\mathbf{x}\in X}\sum_{i=1}^{i_{1}}\lvert u_{i}(\mathbf{x})\rvert^{2}+((i_{2}-i_{1})+\operatorname{S_{\text{tail}}})\max_{i_{1}<i\leq n}\|u_{i}\|_{L^{\infty}(X)}^{2}.$
$\sum_{i=1}^{i_{1}}\lvert u_{i}(\mathbf{x})\rvert^{2}=\|w(\mathbf{x})VD^{-1}\|_{2}^{2}\leq\|w(\mathbf{x})\|_{2}^{2}\|V\|_{2}^{2}\|D^{-1}\|_{2}^{2}=\frac{\|w(\mathbf{x})\|_{2}^{2}}{\sigma_{i_{1}}^{2}}\leq\frac{\|w(\mathbf{x})\|_{2}^{2}}{1-\epsilon^{2}},$
where $w(\mathbf{x})=\begin{bmatrix}\phi_{1,n}(\mathbf{x})&\dots&\phi_{n,n}(\mathbf{x})\end{bmatrix}$, $V\in\mathbb{C}^{n\times i_{1}}$ is a matrix with orthogonal columns and $D\in\mathbb{C}^{i_{1}\times i_{1}}$ is a diagonal matrix with $(D)_{(i,i)}=\sigma_{i}$. The final result follows from
$\sup_{\mathbf{x}\in X}\frac{\|w(\mathbf{x})\|_{2}^{2}}{1-\epsilon^{2}}\leq\frac{1}{1-\epsilon^{2}}\sup_{\mathbf{x}\in X}\sum_{i=1}^{n}\lvert\phi_{i,n}(\mathbf{x})\rvert^{2}\leq\frac{1}{1-\epsilon^{2}}\sum_{i=1}^{n}\|\phi_{i,n}\|_{L^{\infty}(X)}^{2}\leq\frac{n}{1-\epsilon^{2}}.$
@@
Foremost, it becomes clear that regularization decreases the importance of small singular values and the associated basis functions $u_{i}$. For the effective degrees of freedom, it follows from Theorem 11 that $\operatorname{S_{\text{tail}}}=\sum_{i=i_{2}+1}^{n}\left(\sigma_{i}/\epsilon\right)^{2}$ is an important quantity. This quantity heavily depends on the decay of the singular values. For instance, if they are exponentially decaying with a rate that is independent of $n$, one has $\operatorname{S_{\text{tail}}}\leq C$ for varying $n$. In this case, the effective degrees of freedom scale with $i_{2}$. In practice, the decay of the singular values typically has a minor dependence on $n$, see e.g. section 5.2.1, such that $\operatorname{S_{\text{tail}}}$ grows slowly with $n$. Furthermore, we find that the basis functions $u_{i}$ associated with small singular values indeed have a smaller impact on (54) and (55) than those associated to large singular values. This effect can be considerable, as illustrated in section 5.2.2 for the Fourier extension frame, where $k_{n}=\mathcal{O}(n^{2})$ while $k_{n}^{\epsilon}=\mathcal{O}(n\log(n))$.
@@
The reasoning above could give the impression that the behaviour of the basis functions $u_{i}$ associated to small singular values ($i>i_{2}$) is insignificant. According to (54) this is not generally true; only if these functions are sufficiently bounded, the inverse effective Christoffel function $k_{n}^{\epsilon}$ is guaranteed to behave like the inverse Christoffel function $k^{\text{trunc}}_{n}$ associated to the set $\{u_{i}\}_{i=1}^{i_{2}}$. In the discrete setting, the latter is often referred to as the leverage scores associated to the best rank-$i_{2}$ space [9]. This distinction is important for practical purposes, since $k^{\text{trunc}}_{n}$ is computable to high accuracy using $\operatorname{\epsilon_{\text{mach}}}$ precision, while computing $k^{\epsilon}_{n}$ and $k_{n}$ requires extended precision.

### v2

@@ 4 Error analysis of regularized approximations
In the previous sections, we introduced numerical redundancy and identified the numerically achievable accuracy, which differs from that of the best approximation. We also showed that numerical approximations can be interpreted as $\ell^{2}$-regularized approximations. In this section, we examine how regularization influences the errors involved in function approximation. We begin by analysing regularized least squares approximations with a more general penalty term, as our methodology naturally extends to this setting. In section 4.2, we then focus on $\ell^{2}$-regularization, particularly in the context of function approximation using a numerically redundant spanning set.
4.1 General penalty term
@@
with associated operator $\mathcal{F}_{n,m}^{\text{RLS}}$ defined by (7). We assume that $\mathcal{L}_{n,m}^{\text{RLS}}$ is well-defined, i.e., that the minimizer is unique. The penalty function $p_{\epsilon}:\mathbb{C}^{n^{\prime}}\to\mathbb{R}^{+}$ can depend on a regularization parameter $\epsilon$. Furthermore, we assume that $p_{\epsilon}$ is even and that $\sqrt{p_{\epsilon}}$ is subadditive, meaning that for all $\mathbf{x},\mathbf{y}\in\mathbb{C}^{n^{\prime}}$ it holds that
@@
[…]
@@
4.2 Numerical redundancy and $\boldsymbol{\ell^{2}}$-regularization
We are specifically interested in $\ell^{2}$-regularization in order to compute accurate approximations in a numerically redundant spanning set. In this case, one has $p_{\epsilon}(\mathbf{x})=\epsilon^{2}\|\mathbf{x}\|_{2}^{2}$, such that the discretization condition (33) becomes
@@ 4.2.2 Interpretation of the discretization condition
and, hence, for these sets of expansion coefficients (35) is automatically satisfied for $A^{\epsilon}_{n,m}=1$. More generally, (35) is automatically satisfied for any $\mathbf{x}\in\mathbb{C}^{n^{\prime}}$ satisfying $\|\mathcal{T}_{n}\mathbf{x}\|_{H}\leq\epsilon\|\mathbf{x}\|_{2}$. This reflects the idea that regularization makes the effective approximation space smaller.
What happens in the case of an orthonormal spanning set? Such a set satisfies Parseval’s identity: $\|\mathcal{T}_{n}\mathbf{x}\|_{H}=\|\mathbf{x}\|_{2}$. Therefore, (35) is equivalent to
@@
To show how the penalty term affects the discretization condition in a simple setting, consider a basis $\Phi_{n}$ with sampling operator $\mathcal{M}_{m}$, and suppose we add a new function $\psi$ to the basis. Intuitively, increasing the approximation space would require more samples to preserve stability. However, if $\psi$ can be approximated to machine precision by $\Phi_{n}$, then the space hasn’t really grown numerically. As Boyd discusses in [17, Section 16.6], such additions are effectively invisible on a computer. In this case, no extra samples are needed. This can be substantiated using (35), which accounts for finite precision, while the standard condition (9) in general no longer holds (i.e., $A_{n,m}=0$ for the augmented basis).
@@ 4.2.3 Common methods for $\boldsymbol{\ell^{2}}$-regularization
For completeness, we show that TSVD regularization indeed satisfies (40). The same result follows from combining [22, Lemma 3.3] and [8, Theorem 3.8].
@@ 5 Random sampling for $\boldsymbol{L^{2}}$-approximations
The question naturally arises: how can one construct a sampling operator that satisfies the regularized norm inequality (35)? To address this question, we draw inspiration from the unregularized case. Recently, Cohen and Migliorati [21] introduced an elegant strategy to generate random pointwise samples that satisfy (9) for least squares fitting. This approach, reviewed in section 5.1, is closely related to the concept of leverage scores, which are used to subsample tall-and-skinny matrices in a discrete setting. Leverage score sampling is a well-established technique in numerical linear algebra and machine learning. Moreover, an extension of leverage scores, known as ridge leverage scores, has been developed specifically for regularized settings. Using a continuous analogue of these scores, we present a sampling strategy for $\ell^{2}$-regularized function approximation in section 5.2. In section 5.3, a general framework is given for the analysis of the effects of regularization given a numerically redundant spanning set. Throughout the whole section, we work under the mild assumption that for any x in the domain of $V_{n}$, there exists a $v\in V_{n}$ such that $v(\mathbf{x})\neq 0$.
@@ 5.1 Results without regularization
where $\preceq$ denotes the Loewner order. Here, $G_{n,m}$ and $G_{n}$ are the discrete and continuous Gram matrices representing the operators $(\mathcal{M}_{m}\mathcal{U}_{n})^{*}(\mathcal{M}_{m}\mathcal{U}_{n})$ and $\mathcal{U}_{n}^{*}\mathcal{U}_{n}$, respectively, where $\mathcal{U}_{n}$ is the synthesis operator associated with any set of functions $\Psi_{n}$ satisfying $\operatorname{span}(\Psi_{n})=V_{n}$. We stress that this equivalence holds for any such set $\Psi_{n}$, not only for the chosen spanning set $\Phi_{n}$. However, if one considers an orthonormal set $\Psi_{n}=\{\psi_{i}\}_{i=1}^{n}$, the analysis simplifies significantly. In this case, one has $G_{n}=I$ and $A_{n,m}$ is simply related to the smallest eigenvalue of $G_{n,m}$.
As explained in [21, Proof of Theorem 2.1], suppose one uses a sampling operator $\mathcal{M}_{m}$ based on random pointwise samples
@@
This definition assumes that $\Psi_{n}$ is orthonormal. However, it is independent of the specific choice of the orthonormal set $\Psi_{n}$; the function $k_{n}$ only depends on $V_{n}=\operatorname{span}(\Psi_{n})$ and the probability density $\rho$. The function $k_{n}$ has been analysed in the field of classical analysis long before its introduction in the analysis of discrete least squares problems. In classical analysis, it is known as the inverse of the Christoffel function [44]. Furthermore, its discrete analogue is known as leverage scores, which are frequently used in randomized numerical linear algebra and machine learning, see e.g. [41, §2].
This reasoning led to the following sampling result, which is a restatement of [25, Lemma 2.1].
@@
[25, Lemma 2.1] Consider the sampling operator $\mathcal{M}_{m}$ defined by the sampling functionals (43). Let $\gamma>0$ and
@@
It follows that the sampling distribution (46) requires only $\mathcal{O}(n\log(n))$ samples for accurate least squares fitting with high probability. As a result, the amount of required data is nearly optimal, where “optimal” corresponds to the case of interpolation with $m=n$. For a summary of optimality benchmarks and probabilistic error analyses, we refer to [25, sections 1 and 2]. Also, we note that considerable efforts have been made to obtain $\mathcal{O}(n)$ sampling strategies, see e.g. [1, section 8] and references therein. However, these methods are no longer purely random, which diverges from the focus of this section.
5.2 Results with $\boldsymbol{\ell^{2}}$-regularization
@@
An important difference compared to section 5.1 is the dependence on the chosen spanning set $\Phi_{n}$. In this case, we are bound to the choice $\Psi_{n}=\Phi_{n}$, i.e., $G_{n,m}$ and $G_{n}$ represent the operators $(\mathcal{M}_{m}\mathcal{T}_{n})^{*}(\mathcal{M}_{m}\mathcal{T}_{n})$ and $\mathcal{T}_{n}^{*}\mathcal{T}_{n}$, respectively, where $\mathcal{T}_{n}$ is the synthesis operator associated with the spanning set $\Phi_{n}$.
Similar spectral bounds have been analysed before in the setting of randomized algorithms [11, 12, 13, 48] using analogous techniques as the ones described in section 5.1. In this case, a continuous analogue of the $\epsilon$-ridge leverage scores [11, Definition 1] is of interest
@@
where $\sigma_{i}$ and $u_{i}(\mathbf{x})$ are the singular values and left singular vectors, evaluated at $\mathbf{x}$, associated with the synthesis operator $\mathcal{T}_{n}$. We refer to (48) as the inverse of the numerical Christoffel function.
@@
where we call the quantity $n^{\epsilon}$ the numerical dimension. In the discrete context, $n^{\epsilon}$ is also known as the effective dimension or the effective degrees of freedom [11]. It satisfies $n^{\epsilon}\leq n$ and it is a decreasing function of the regularization parameter $\epsilon$.
Using very similar techniques as [12, proof of Lemma 6], we can prove the following sampling result. The reworked proof can be found in Appendix B. [^2] We believe a result involving the $L^{\infty}(X)$-norm might be true, as in Theorem 9. However, the proof of Theorem 10 relies on the matrix concentration inequality given in [12, Lemma 27] which requires the use of the uniform norm.
@@
$m\geq\frac{32}{3}\|wk_{n}^{\epsilon}\|_{\infty}\log(16n^{\epsilon}/\gamma),$
where $\|wk_{n}^{\epsilon}\|_{\infty}=\sup_{\mathbf{x}\in X}\lvert w(\mathbf{x})k_{n}^{\epsilon}(\mathbf{x})\rvert$, then (35) is satisfied for $A_{n,m}^{\epsilon}=1/2$ with probability at least $1-\gamma$.
@@
It is worthwhile to interpret the differences between the results laid out in section 5.1 and those in this section. Using Theorem 1 the sampling strategy proposed in Corollary 1 results in a least squares approximation with
@@
samples, again with high probability and assuming the data $\mathbf{f}$ is noiseless. In these statements the constant $C>0$ is generic and may change with each occurrence. These results demonstrate that regularization generally decreases the achievable accuracy, yet also decreases the amount of data required to obtain a near-best approximation. Furthermore, the numerical dimension $n^{\epsilon}$ in regularized approximations serves a similar role as the dimension $n$ in unregularized approximations.
Note that in practice one typically has $n^{\epsilon}=\mathcal{O}(n)$ such that (52) and (54) are asymptotically equivalent as $n\to\infty$. This suggests that the gain in data efficiency achieved through regularization is generally limited to a constant factor when samples are drawn from the optimal distributions proposed in Corollary 1 and 2. On the other hand, regularization can have a more substantial influence on the required sample complexity if one cannot choose the sample distribution. Since $k_{n}$ can differ significantly from $k_{n}^{\epsilon}$, the required number of samples drawn from a specific distribution $\mu$ may vary considerably depending on whether or not the effects of regularization are accounted for, as described in Theorems 9 and 10. An illustration of this phenomenon are the results of section 6.2.2 for Fourier extension frames.
@@ 5.3 Behaviour of $\boldsymbol{n^{\epsilon}}$ and $\boldsymbol{k_{n}^{\epsilon}}$
For approximation with numerically redundant spanning sets, we are specifically interested in the setting where $\epsilon$ is of the order of a finite working precision $\operatorname{\epsilon_{\text{mach}}}$ associated with the numerical computations. It is important to gain intuition in the impact of the penalty term, i.e., whether the numerical dimension $n^{\epsilon}$ differs significantly from $n$ and the behaviour of $k^{\epsilon}_{n}$ differs significantly from the behaviour of $k_{n}$, for small regularization parameters $\epsilon$.
@@
The region in between $i_{1}$ and $i_{2}$ is often referred to as the plunge region [23]. These regions have a different influence on the numerical dimension $n^{\epsilon}$ and the maximum of the inverse numerical Christoffel function $k_{n}^{\epsilon}$, as quantified in the following theorem, of which the proof can be found in Appendix C.
@@
where $\operatorname{S_{\text{tail}}}=\sum_{i=i_{2}+1}^{n}\left(\sigma_{i}/\epsilon\right)^{2}$ and $k^{\text{trunc}}_{n}(\mathbf{x})=\sum_{i=1}^{i_{2}}\lvert u_{i}(\mathbf{x})\rvert^{2}$. Furthermore, if $\Phi_{n}$ is linearly independent and uniformly bounded, $\|\phi_{i,n}\|_{\infty}\leq 1\;(1\leq i\leq n)$, then
$\|k_{n}^{\epsilon}\|_{\infty}\leq\frac{n}{1-\epsilon^{2}}+((i_{2}-i_{1})+\operatorname{S_{\text{tail}}})\max_{i_{1}<i\leq n}\|u_{i}\|_{\infty}^{2}.$ (59)
@@
Foremost, it becomes clear that regularization decreases the importance of small singular values and the associated basis functions $u_{i}$. For the numerical dimension, it follows from Theorem 11 that $\operatorname{S_{\text{tail}}}=\sum_{i=i_{2}+1}^{n}\left(\sigma_{i}/\epsilon\right)^{2}$ is an important quantity. This quantity heavily depends on the decay of the singular values. For instance, if they are exponentially decaying with a rate that is independent of $n$, one has $\operatorname{S_{\text{tail}}}\leq C$ for varying $n$. In this case, the numerical dimension scales with $i_{2}$. In practice, the decay of the singular values typically has a minor dependence on $n$, see e.g. section 6.2.1, such that $\operatorname{S_{\text{tail}}}$ grows slowly with $n$. Furthermore, we find that the basis functions $u_{i}$ associated with small singular values indeed have a smaller impact on (58) and (59) than those associated to large singular values. This effect can be considerable, as illustrated in section 6.2.2 for the Fourier extension frame, where $k_{n}=\mathcal{O}(n^{2})$ while $k_{n}^{\epsilon}=\mathcal{O}(n\log(n))$.
@@
The reasoning above could give the impression that the behaviour of the basis functions $u_{i}$ associated to small singular values ($i>i_{2}$) is insignificant. According to (58) this is not generally true; only if these functions are sufficiently bounded, is the inverse numerical Christoffel function $k_{n}^{\epsilon}$ guaranteed to behave like the inverse Christoffel function $k^{\text{trunc}}_{n}$ associated to the set $\{u_{i}\}_{i=1}^{i_{2}}$. In the discrete setting, the latter is referred to as the leverage scores associated to the best rank-$i_{2}$ space [11]. This distinction is important for practical purposes, since $k^{\text{trunc}}_{n}$ is computable to high accuracy using $\operatorname{\epsilon_{\text{mach}}}$ precision, while computing $k^{\epsilon}_{n}$ and $k_{n}$ requires extended precision.